OpenStudy (anonymous):

Is this correct? 2(3^-2)(-2^-2) = 1/2(3^2)(-2^2) = 1/72 ?

OpenStudy (anonymous):

um its 2*1/9*-1/4=-1/18 :)

OpenStudy (campbell_st):

I think its wrong \[2(3^{-2})(-2^{-2}) = 2/(3^2)(-2^2) = 1/18\]

OpenStudy (campbell_st):

remember, square a negative gives a positive

OpenStudy (anonymous):

i think you just multipled wrong, youe have the correct idea. also remember this when you have -x^2 and your told x=1 this does not mean (-1)^2=1 it means -(1)^2=-1 :)

OpenStudy (anonymous):

nooo campbell_st you are wrong, unless she meant to put brackets it is wrong

OpenStudy (anonymous):

so it's positive 18 or -18 ?

OpenStudy (campbell_st):

then if I'm wrong the answer will be \[2/(3^2)-(2)^2 = 2/5\]

OpenStudy (campbell_st):

its positive 18...

OpenStudy (anonymous):

are there brackets around the -2^-2? is it technicaly like this -(2)^-2 or like this (-2)^-2?

OpenStudy (campbell_st):

so your telling me (-2 x -2 ) = -4

OpenStudy (anonymous):

you are missing brackets campbell...

OpenStudy (anonymous):

ur not understanding

OpenStudy (anonymous):

i am saying this -(2)^2= -1*(2*2)=-4 do you understand?

OpenStudy (campbell_st):

then the answer is 9 - 4 = 5 so its 2/5

OpenStudy (anonymous):

justine the way your question is written it should be negative

OpenStudy (anonymous):

WHY ARE YOU SUBTRACTING?!

OpenStudy (campbell_st):

I see the neagtive and square inside the brackets... and squaring a negative gives a positive you keep telling me the negative is an operator....

OpenStudy (anonymous):

LOOK IF IT IS WRITTEN AS SUCH -(x)^2 then the answer will be - BUT if it is as you interpreted it (-x)^2 then the answer will b + get it??

OpenStudy (campbell_st):

well I hope you have a lot of success with that method

OpenStudy (anonymous):

campbell, you want MORE proof? http://www.wolframalpha.com/input/?i=2%283%5E-2%29%28-2%5E-2%29

OpenStudy (anonymous):

i just put in her exact equation look at the answer

OpenStudy (anonymous):

math is not like english or french, etc its simple. there is one way of speaking one correct answer. -1/18 is the correct answer jasminee sorry for all this but I hope you understand because it is very easy to make mistakes like this

OpenStudy (anonymous):

campbell you arrogance will not do you well, be open to learning new things. I have made this mistake before. Important thing is to learn from them. I have nothing against you, i am sure you are brilliant. but just be a little more open thats all.